OpenStudy (anonymous):

Alan leaves Los Angeles at 8:00 a.m. to drive to San Francisco, 400mi\;{\rm mi} away. He travels at a steady 50mph\;{\rm mph} . Beth leaves Los Angeles at 9:00 a.m. and drives a steady 60mph\;{\rm mph} .B How long does the first to arrive have to wait for the second?

OpenStudy (anonymous):

help

OpenStudy (xapproachesinfinity):

\;{\rm mi} what is this?

OpenStudy (anonymous):

I don't know it to hard of a question

OpenStudy (xapproachesinfinity):

can you do a snapshot of the problem

OpenStudy (anonymous):

oh yeah it just mph

OpenStudy (anonymous):

no rm

OpenStudy (anonymous):

it just mi and mph

OpenStudy (anonymous):

do you get it

OpenStudy (anonymous):

can get the screenshot dude

OpenStudy (xapproachesinfinity):

okay so! we need to know how many our each will drive to san francisco the equation is like this v=d/t so t=d/v we are given the distance btw los and san to be 400 v=50 for alan so for Alan we have t=400/50=8 hour so if he started at 8:00 am he will arrive at 4:00pm

OpenStudy (anonymous):

bent is the one that arrive first

OpenStudy (xapproachesinfinity):

for beth t=400/60=6.66 hour that is 6hour 40min so if he started at 9:00am he will arrive at 3:40 pm

OpenStudy (xapproachesinfinity):

yeah so he will arrive 20 minutes before Alan

OpenStudy (xapproachesinfinity):

Do you get the process here?

OpenStudy (anonymous):

yeah but how much times the another one have to what for the other one

OpenStudy (anonymous):

is it 20 min

OpenStudy (xapproachesinfinity):

yes 20 minute since the first arrives at 3.40 the second 4:00 so 20 minute difference

OpenStudy (anonymous):

can you help my on a another one just one more